# Physics of droplet regulation in biological cells

### v2

@@ 4.3 Interaction with membranes
Cellular membranes take the form of phospholipid bilayers, and allow the cell to partition different regions of space. Membranes are typically very thin and flexible, and contain an array of proteins and other biomolecules that can reorganise laterally in a fluid-like manner. Droplets come into contact with many different membranes within the cell, including the plasma membrane, membrane-bound organelles, vesicles, and the endoplasmic reticulum. Droplet–membrane interactions are mediated by both physical (e.g., wetting) and biomolecular (e.g., binding) effects [349, 350, 351, 60]. In this section, we first discuss interactions between macroscopic droplets and membranes, before considering the possibility of phase separation phenomena within the membrane itself.
@@ 4.3.1  Condensation at membranes
One striking example of membrane–droplet interactions is the complex deformation and mutual remodelling that results from wetting effects, as highlighted by recent experiments with synthetic analogue systems [351, 60] and detailed numerical simulations [359]. This two-way coupling causes membrane budding, interfacial ruffling, and nanotube formation. As discussed in section 4.1.2 above, capillary forces are able to deform soft surfaces over the scale of the elasto-capillary length $\ell_{\textrm{cap}}$. For deformable solids which are extremely thin in one dimension, such as membranes, the relevant elastic energy which resists capillary forces is the bending energy, characterised by bending rigidity $B$. Redefining the elasto-capillary length as $\ell_{\textrm{cap}}=\gamma/B$ then gives the length scale over which capillary forces give rise to membrane–droplet deformations.
@@ 4.4.1  Network-excluding droplets
where $\gamma$ is surface tension, and $\Delta\!f$ is the driving force behind droplet formation; see equation (2.33). The elastic pressure in the surrounding mesh is given by $P_{\mathrm{el}}=-\partial F_{\textrm{el}}/\partial V$, and is a function of droplet size. The specific form of $P_{\mathrm{el}}(R)$ depends on the particular rheological response of the mesh. Minimising equation (4.9), we see that elastic stresses can stabilise droplets, with a stable droplet radius $R_{*}$ given by
@@
To capture the dynamics of droplets embedded in elastic meshes, we describe the dilute and dense phases separately, as in section 2.2. If multiple droplets exist, then the evolution of droplet sizes is given by equation (2.44), but with the equilibrium volume fraction of the dilute phase, $\phi_{\textrm{out}}^{\textrm{eq}}$, now a function of $P_{\textrm{el}}$, which in turn depends on the size of the droplet. In general, $P_{\textrm{el}}$ can also depend on space, as discussed in section 4.4.3 below. Note that the presence of the mesh restricts droplet motion, such that droplets can be assumed to be fixed in place once they have nucleated. The evolution of the dilute phase follows equation (2.43).
@@ 4.4.2  Network permeation
where $\ell_{\mathrm{p}}$ is the characteristic pore size of the network. This wetting energy represents an averaging of capillary interactions between the droplet and individual filaments over a representative volume element of the mesh. A simple scaling analysis shows that the transition between network permeation and network exclusion is controlled by the ‘permeo-elastic number’ $\mathcal{P}=(\gamma_{\mathrm{in}}-\gamma_{\mathrm{out}})/(\ell_{\mathrm{p}}E)$, for elastic modulus $E$.
@@ 4.4.3  Complex properties of the mesh
So far, we have focussed on homogeneous meshes; in reality, meshes within cells display both small-scale and large-scale heterogeneity [383]. When droplets are the same size as characteristic mesh heterogeneities, bulk elasticity is not a good representation of the mesh’s mechanical response. Instead, non-local effects from neighbouring mesh filaments must be taken into account. Non-local elasticity can be included in the current framework by including an elastic contribution to the free energy given by a convolution of the elastic pressure in the mesh with a representative non-locality kernel [384]. Non-local elasticity has been shown to facilitate the formation of equilibrium patterned phases during phase separation [384, 385], as observed in recent experiments with synthetic polymer meshes [386].
@@
Many biological meshes, such as the cytoskeleton, also display viscoelastic behaviour over sufficiently long timescales. As a result, elastic stresses can relax over time, facilitating further droplet growth, which may in turn lead to accelerated ripening [388]. Finally, although we have focussed on passive meshes here, molecular motors may also drive active contraction of the cytoskeleton [389], in turn deforming droplets that bind to the mesh [390]. The interaction between biomolecular condensates and such active meshes remains relatively unexplored.
@@ 5.1 Chemical reactions of many components
We start by considering a homogeneous mixture of ${N_{\mathrm{c}}}$ components $X_{i}$ that participate in ${N_{\mathrm{r}}}$ reactions,
@@
where $\sigma^{(\alpha)}_{\rightarrow,i}$ and $\sigma^{(\alpha)}_{\leftarrow,i}$ are the respective stoichiometry coefficients of reactants and products for ${\alpha=1,\ldots,{N_{\mathrm{r}}}}$. The system is in equilibrium when the chemical potentials $\mu_{i}$ are balanced for each reaction,
@@
where the square bracket describes energetic constraints, whereas kinetic effects are captured by the positive rate coefficient $k_{\alpha}$, which can depend on the composition $\Phi$. The functional form of $k_{\alpha}$ can be constrained by requiring that $s^{(\alpha)}$ reduces to mass action kinetics for ideal solutions, implying that (appendix F)
@@ 5.2 Chemically active binary mixtures
where $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are the respective reaction coefficients for the passive and active reactions. Both $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are positive and may depend on the composition $\phi$.
@@ 5.2.1  Stability of homogeneous states
To build intuition, we first analyse the stability of homogeneous stationary states, which obey $\phi(\boldsymbol{r})=\phi_{0}$ with $s(\phi_{0})=0$. Considering harmonic perturbations with wave vector $\boldsymbol{q}$, the associated growth rate $\omega$ is equal to the single entry of the Hessian given by equation (2.15),
@@ 5.2.2  Effective droplet model
To understand the dynamics of chemically active droplets in more detail, we next discuss the effective droplet model introduced in section 2.2. The equilibration of the interfacial region is not affected significantly by reactions since the interfacial volume is typically small. Instead, reactions affect the dynamical equation in the bulk phases, which after linearisation led to the diffusion equation (2.37). Repeating this linearisation for equation (5.7), we find
$\partial_{t}\phi_{n}\approx D_{n}\nabla^{2}\phi_{n}+\Gamma_{n}-k_{n}(\phi-\phi^{(0)}_{n})\;,$ (5.9)
@@ 5.5.2  Self-organized drift
Chemical gradients also induce surface tension gradients, which can drive movement via Marangoni flows, as discussed in section 3.3. While droplets generically react to external gradients, there is also the exciting possibility that they may create their own gradients. Moreover, they may exploit an instability where movement amplifies this gradient, so that the entire system settles into a moving state, either via diffusiophoresis [411] or Marangoni flows [272, 64]. Alternatively, chemical reactions can create an asymmetric state, which then moves due to isotropies in surrounding gradients [179]. In all these cases, activity needs to exceed a threshold, above which isotropy is spontaneously broken and swimming begins.
@@ 5.5.3  Active multi-component mixtures
where $s^{(\alpha)}$ denotes the flux of reaction $\alpha$, e.g., given by equation (5.4), and we neglect hydrodynamic effects and thermal noise for simplicity. These equations describe how chemical reactions affect droplets and how droplets affect chemical reactions [442].
Active multi-component mixtures can exhibit various instabilities and patterns, ranging from simple Turing patterns [443, 402, 403, 429] to systems with multiple length scales [444], to more complex spatiotemporal behaviour [445, 446]. Particularly interesting is the case where phase separation interferes with chemical reactions [447, 448], so that behaviours beyond those described above can emerge. For instance, droplets may now be stable even in the presence of passive reactions, which is impossible in binary systems [399]. In this case, internally maintained droplets do not fill the entire system, although their size typically still scales with system size [408, 392]. Externally maintained droplets, however, can either scale with system size or exhibit the size control discussed in section 5.3 [415, 449]. In both of these cases, the droplet size distribution relaxes to a monodisperse (and universal) distribution at late times [449]. Externally maintained droplets can also assume shell-like structures [410, 450], essentially because chemical reactions deplete the inside of the droplet so that concentration (see figure 5.3a) drops below the spinodal concentration and a bubble emerges. Chemically active droplets may be further controlled by enzyme positioning, either because they are mobile [451] or they are fixed in particular locations [452]. Recent work has also shown that enzymes themselves may form phase-separated clusters with a characteristic equilibrium size [453], similar to the size control exhibited by externally maintained droplets discussed in section 5.3.1. As this field becomes more developed, there will surely be many additional phenomena that we are yet unveil.
@@ 6.1 Life cycle of droplets in biological cells
In many situations, cells need to control when and where droplets form. Such control is difficult to achieve via classical spontaneous nucleation. Although heterogeneous nucleation sites can bias droplet position (section 4.1.4), true spatial control is best realised by the use of active catalytic sites (section 5.4), which drive localised droplet formation in an undersaturated system. The timing of droplet formation can be controlled by actively driven bulk reactions, which can suppress phase separation (section 5.2.1). This process requires a continuous expenditure of fuel such as ATP, but has the advantage that droplets form immediately when fuel ceases, thus providing a robust stress response [454].
@@
Droplets need to be in the right place to fulfil their function. To control position, cells can anchor droplets at particular structures, such as the plasma membrane or chromosome loci, so that they remain in place. If droplets are not anchored, cells can reposition them, e.g., using external gradients (section 2.3.2) and hydrodynamic fluxes. Even more exciting are droplets that swim themselves, either because they self-organise chemical gradients (section 5.5.2) or they exert active stresses [84, 460, 461].
@@ 6.2 Droplet–environment interactions
The interplay of droplets with each other and their environment leads to complex interaction networks, which can process information in cells and enable communication between droplets [467]. Such interaction networks typically involve multiple feedback loops, which give rise to additional dynamic behaviours [468]. At an abstract level, droplets might thus play an integral part in information processing and learning in biological cells [318].
@@ 6.3 Outlook
Beyond exploring how droplets enable cellular function, understanding the physical principles that biology exploits so effectively has application to other fascinating fields of research. For example, it has been hypothesised that phase separation could be implicated in the origin of life by enabling the formation of basic protocells [139]. Indeed, droplets provide well-defined entities that grow, move, and divide by metabolising material from their environment [470]. On the more theoretical side, chemically active droplets comprise an exciting class of non-equilibrium systems [84], with similarities to proliferating active matter [471] and general dissipative structures [472]. Finally, understanding droplet regulation in biological cells will allow us to leverage similar principles in designing novel synthetic systems [61, 66].
We would like to thank Jan Kirschbaum and Estefania Vidal-Henriquez for early input and drafting initial concepts for the review. We thank Stefan Köstler, Guido Kusters, Henri Schmidt, Filipe Thewes, and Noah Ziethen for helpful comments and discussions of the manuscript. We further thank Kristian Blom, Gerrit Wellecke, Riccardo Rossetto, Yicheng Qiang, Marcel Ernst, Chengjie Luo, and the entire Zwicker group for invaluable scientific discussions. We gratefully acknowledge funding from the Max Planck Society and the European Union (ERC, EmulSim, 101044662).
@@ A Compressible multi-component fluid
The particle fluxes $\boldsymbol{\hat{J}}_{i}$ can be split into a centre-of-mass motion and relative particle fluxes ${\boldsymbol{\hat{j}}_{i}=\boldsymbol{\hat{J}}_{i}-c_{i}\boldsymbol{v}}$ [96], implying the constraint $\sum_{i=0}^{{N_{\mathrm{c}}}}m_{i}\boldsymbol{\hat{j}}_{i}=0$. The dynamics of the flow field $\boldsymbol{v}$ follow from momentum conservation, $\partial_{t}(\rho\boldsymbol{v})=\nabla\cdot\boldsymbol{\sigma}^{\mathrm{tot}}$, where the total stress $\boldsymbol{\sigma}^{\mathrm{tot}}=\boldsymbol{\sigma}-\rho\boldsymbol{v}\boldsymbol{v}$ depends on the stress $\boldsymbol{\sigma}$ in the co-moving reference frame [115]. The relative fluxes $\boldsymbol{\hat{j}}_{i}$ can be expanded assuming a linear response to thermodynamic forces [96],
@@
where we have introduced the exchange chemical potentials $\hat{\bar{\mu}}_{j}=\hat{\mu}_{j}-m_{j}m_{0}^{-1}\hat{\mu}_{0}$, since only ${N_{\mathrm{c}}}$ fluxes are independent and contribute to entropy production, whereas $\boldsymbol{\hat{j}}_{0}=-m_{0}^{-1}\sum_{i=1}^{N_{\mathrm{c}}}m_{i}\boldsymbol{\hat{j}}_{i}$. The noise $\boldsymbol{\hat{\xi}}_{i}$ obeys the fluctuation–dissipation relation [96]
$\langle\boldsymbol{\hat{\xi}}_{i}(\boldsymbol{r},t)\boldsymbol{\hat{\xi}}_{j}(\boldsymbol{r}^{\prime},t^{\prime})\rangle=2k_{\mathrm{B}}T\hat{\Lambda}_{ij}\mathbb{1}\delta(\boldsymbol{r}-\boldsymbol{r}^{\prime})\delta(t-t^{\prime})\;.$ (A.6)
The ${N_{\mathrm{c}}}$-dimensional mobility matrix $\hat{\Lambda}_{ij}$ obeys Onsager’s symmetry condition, $\hat{\Lambda}_{ij}=\hat{\Lambda}_{ji}$. Note that we chose to express fluxes and chemical potentials relative to the solvent ($i=0$), but other choices are possible. To reflect this symmetry, we introduce the $({N_{\mathrm{c}}}+1)$-dimensional mobility matrix $\hat{L}_{ij}$, which captures the mobilities of all individual components. The two mobility matrices are linked by
@@
implying that $\sum_{i=0}^{N_{\mathrm{c}}}m_{i}\hat{\Lambda}_{ij}=0$.
To describe liquids, we take the incompressible limit ($K\rightarrow\infty$), which enforces a constant volume density, $\sum_{i}\nu_{i}c_{i}=1$. In the simple case where all components have equal mass density $m_{i}/\nu_{i}$, the total mass density $\rho$ is then also conserved, implying incompressible flows (.$\nabla.\boldsymbol{v}=0$) enforced by hydrostatic pressure. The system state is then conveniently expressed in terms of volume fractions $\phi_{i}=\nu_{i}c_{i}$, with the solvent fraction $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$. The kinetics described by equations (2.8)–(2.11) follow directly by substituting $\Lambda_{ij}=\nu_{i}\hat{\Lambda}_{ij}$ and $\boldsymbol{j}_{i}=\nu_{i}\boldsymbol{\hat{j}_{i}}$. The associated mobility matrix given by equation (2.12) follows directly from equation (A.7). In the simple case of a non-interacting system ($h=0$), diagonal mobilities, $L_{ij}=\lambda_{i}\phi_{i}\delta_{ij}$, dilute concentrations ($\phi_{i}\ll 1$ for $i\geq 1$), and a fast solvent ($\lambda_{0}\gg\lambda_{i}$), the kinetics reduce to ideal diffusion, $\boldsymbol{j}_{i}=-D_{i}\nabla\phi_{i}$, with diffusivity $D_{i}=k_{\mathrm{B}}T\lambda_{i}$.
@@
where $e_{i}$ are internal energies and $e_{ij}$ denote interaction energies obeying $e_{ij}=e_{ji}$. This model can be derived from a lattice model or can be viewed as a simple polynomial expansion up to second order, which is sufficient to describe phase separation, although higher-order effects are also interesting [321]. Taken together, this results in the incompressible free energy density given in equation (2.1), where
@@
with $\chi_{ij}=\chi_{ji}$, but $\chi_{ii}\neq 0$. The exchange chemical potentials in equation (2.2) follow from equation (A.3) using $\bar{\mu}_{i}=\hat{\mu}_{i}-\Omega_{i}\hat{\mu}_{0}$ with $\Omega_{i}=\nu_{i}/\nu_{0}$. Similarly, the pressure $P=\hat{P}-\hat{\mu}_{0}\nu_{0}^{-1}$ in equation (2.3) follows from equation (A.4), illustrating that the osmotic pressure $P$ involves the chemical potential of the solvent.

### v3

@@ 4.3 Interaction with membranes
Cellular membranes take the form of phospholipid bilayers, and allow the cell to partition different regions of space. Membranes are typically very thin and flexible, and contain an array of proteins and other biomolecules that can reorganise laterally in a fluid-like manner. Droplets come into contact with many different membranes within the cell, including the plasma membrane, membrane-bound organelles, vesicles, and the endoplasmic reticulum. Droplet–membrane interactions are mediated by both physical (e.g., wetting) and biomolecular (e.g., binding) effects [386, 387, 388, 64, 389, 390]. In this section, we first discuss interactions between macroscopic droplets and membranes, before considering the possibility of phase separation phenomena within the membrane itself.
@@ 4.3.1 Condensation at membranes
One striking example of membrane–droplet interactions is the complex deformation and mutual remodelling that results from wetting effects, as highlighted by recent experiments with synthetic analogue systems [388, 64, 389, 398] and detailed numerical simulations [399]. This two-way coupling causes membrane budding, interfacial ruffling, and nanotube formation. As discussed in section 4.1.2 above, capillary forces are able to deform soft surfaces over the scale of the elasto-capillary length $\ell_{\textrm{cap}}$. For deformable solids which are extremely thin in one dimension, such as membranes, the relevant elastic energy which resists capillary forces is the bending energy, characterised by bending rigidity $B$. Redefining the elasto-capillary length as $\ell_{\textrm{cap}}=\gamma/B$ then gives the length scale over which capillary forces give rise to membrane–droplet deformations.
@@ 4.4.1 Network-excluding droplets
where $\gamma$ is surface tension, and $\Delta\!f$ is the driving force behind droplet formation; see equation (2.33). The elastic pressure in the surrounding mesh is given by $P_{\mathrm{el}}=-\partial F_{\textrm{el}}/\partial V$, and is a function of droplet size. The specific form of $P_{\mathrm{el}}(R)$ depends on the particular rheological response of the mesh. For a droplet growing in a neo-Hookean mesh with elastic modulus $E$, $P_{\mathrm{el}}$ is a monotonically increasing function of $R$ that converges to $P_{\mathrm{el}}=5E/6$ at large $R$ [141]. Minimising equation (4.8), we see that elastic stresses can stabilise droplets, with a stable droplet radius $R_{*}$ given by the implicit equation
@@
To capture the dynamics of droplets embedded in elastic meshes, we describe the dilute and dense phases separately, as in section 2.2. If multiple droplets exist, then the evolution of droplet sizes is given by equation (2.44), but with the equilibrium volume fraction of the dilute phase, $\phi_{\textrm{out}}^{\textrm{eq}}$, now a function of $P_{\textrm{el}}$, which in turn depends on the size of the droplet. The dependence of $\phi_{\textrm{out}}^{\textrm{eq}}$ on $P_{\textrm{el}}$ can be found by considering a pressure balance between the dilute and dense phases [144]. In general, $P_{\textrm{el}}$ can also depend on space, as discussed in section 4.4.3 below. Note that the presence of the mesh restricts droplet motion, such that droplets can be assumed to be fixed in place once they have nucleated. The evolution of the dilute phase follows equation (2.43).
@@ 4.4.2 Network permeation
where $\ell_{\mathrm{p}}$ is the characteristic pore size of the network, and we recall that $\gamma_{\mathrm{in}/\mathrm{out}}$ are the surface energies corresponding to network–droplet and network–solvent interfaces, respectively. This wetting energy represents an averaging of capillary interactions between the droplet and individual filaments over a representative volume element of the mesh. A simple scaling analysis shows that the transition between network permeation and network exclusion is controlled by the ‘permeo-elastic number’ $\mathcal{P}=(\gamma_{\mathrm{in}}-\gamma_{\mathrm{out}})/(\ell_{\mathrm{p}}E)$, for elastic modulus $E$.
@@ 4.4.3 Complex properties of the mesh
So far, we have focussed on homogeneous meshes; in reality, meshes within cells display both small-scale and large-scale heterogeneity [423]. When droplets are the same size as characteristic mesh heterogeneities, bulk elasticity is not a good representation of the mesh’s mechanical response. Instead, non-local effects from neighbouring mesh filaments must be taken into account. Non-local elasticity can be included in the current framework by including an elastic contribution to the free energy given by a convolution of the elastic pressure in the mesh with a representative non-locality kernel [424]. Non-local elasticity has been shown to facilitate the formation of equilibrium patterned phases during phase separation [424, 425, 426, 427], as observed in recent experiments with synthetic polymer meshes [428].
@@
Many biological meshes, such as the cytoskeleton, also display viscoelastic behaviour over sufficiently long timescales. As a result, elastic stresses can relax over time, facilitating further droplet growth, which may in turn lead to accelerated ripening [430]. Finally, although we have focussed on passive meshes here, molecular motors may also drive active contraction of the cytoskeleton [431], in turn deforming droplets that bind to the mesh [432], and potentially leading to droplet size selection [433]. The interaction between biomolecular condensates and such active meshes remains relatively unexplored.
@@ 5.1 Chemical reactions of many components
We start by considering a homogeneous mixture of $N_{\mathrm{c}}$} components $X_{i}$ that participate in $N_{\mathrm{r}}$} reactions [437],
@@
where $\sigma^{(\alpha)}_{\rightarrow,i}$ and $\sigma^{(\alpha)}_{\leftarrow,i}$ are the respective stoichiometry coefficients of reactants and products for $\alpha=1,\ldots,{N_{\mathrm{r}}}$}. If the system is at thermodynamic equilibrium, the respective free energy is minimized, implying that chemical potentials $\mu_{i}$ are balanced for each reaction,
@@
where the second term describes thermodynamic constraints, whereas kinetic effects are captured by the positive rate coefficient $k_{\alpha}$, which can depend on the composition $\Phi$. The functional form of $k_{\alpha}$ can be constrained by requiring that $s^{(\alpha)}$ reduces to mass action kinetics for ideal solutions, implying that (appendix F)
@@ 5.2 Chemically active binary mixtures
where $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are the respective reaction coefficients for the passive and active reactions. Both $k_{\mathrm{p}}$ and $k_{\mathrm{a}}$ are positive and may depend on the composition $\phi$ and the chemical potential $\bar{\mu}$.
@@ 5.2.1 Stability of homogeneous states
To build intuition, we first analyse the stability of homogeneous stationary states, which obey $\phi(\boldsymbol{r})=\phi_{0}$ with $s(\phi_{0})=0$. Considering harmonic perturbations with wave vector $\boldsymbol{q}$, the associated growth rate $\omega$ is equal to the single entry of the Jacobian given by equation (2.15),
@@ 5.2.2 Effective droplet model
To understand the dynamics of chemically active droplets in more detail, we next discuss the thin-interface model introduced in section 2.2. The equilibration of the interfacial region is not affected significantly by reactions since the interfacial volume is typically small. Instead, reactions affect the dynamical equation in the bulk phases, which after linearisation led to the diffusion equation (2.37). Repeating this linearisation for equation (5.7), we find
$\partial_{t}\phi_{n}\approx D_{n}\nabla^{2}\phi_{n}+\Gamma_{n}-k_{n}(\phi_{n}-\phi^{(0)}_{n})\;$, (5.9)
@@ 5.5.2 Self-organized drift
Chemical gradients also induce surface tension gradients, which can drive movement via Marangoni flows, as discussed in section 3.3. While droplets generically react to external gradients, there is also the exciting possibility that they may create their own gradients. Moreover, they may exploit an instability where movement amplifies this gradient, so that the entire system settles into a moving state, either via diffusiophoresis [454] or Marangoni flows [294, 68]. Alternatively, chemical reactions can create an asymmetric state, which then moves due to isotropies in surrounding gradients [189], and even induce active capillary waves [485]. In all these cases, activity needs to exceed a threshold, above which isotropy is spontaneously broken and swimming begins. In contrast, isotropy can also be broken in multiphasic systems, which can exhibit self-propulsion at arbitrarily low activity [486].
@@ 5.5.3 Active multi-component mixtures
where $s^{(\alpha)}$ denotes the flux of reaction $\alpha$, e.g., given by equation (5.4), and we neglect hydrodynamic effects and thermal noise for simplicity. These equations capture how chemical reactions affect droplets and how droplets affect chemical reactions [487].
Active multi-component mixtures can exhibit various instabilities and patterns, ranging from simple Turing patterns [488, 445, 446, 472] to systems with multiple length scales [489], to more complex spatiotemporal behaviour [490, 491, 492]. Particularly interesting is the case where phase separation interferes with chemical reactions [493, 494], so that behaviours beyond those described above can emerge. For instance, droplets may now be stable even in the presence of passive reactions, which is impossible in binary systems [442]. In this case, internally maintained droplets do not fill the entire system, although their size typically still scales with system size [451, 435]. Externally maintained droplets, however, can either scale with system size or exhibit the size control discussed in section 5.3 [458, 495]. In both of these cases, the droplet size distribution relaxes to a monodisperse (and universal) distribution at late times [495]. Externally maintained droplets can also assume shell-like structures [453, 496], essentially because chemical reactions deplete the inside of the droplet so that concentration (see figure 5.3a) drops below the spinodal concentration and a bubble emerges. Chemically active droplets may be further controlled by enzyme positioning, either because they are mobile [497] or they are fixed in particular locations [498]. Recent work has also shown that enzymes themselves may form phase-separated clusters with a characteristic equilibrium size [499], similar to the size control exhibited by externally maintained droplets discussed in section 5.3.1. As this field becomes more developed, there will surely be many additional phenomena that we are yet to unveil.
We here presented a continuous theory of chemically active droplets, which negelects a lot of molecular details of the involved biomolecules. For instance, even simple molecular effects, like crowding, can affect active droplet [500]. In the future, it will thus be important to study thermodynamically consistent molecular models of chemically active droplets [501].
@@ 6.1 Life cycle of droplets in biological cells
In many situations, cells need to control when and where droplets form. Such control is difficult to achieve via classical spontaneous nucleation (section 2.2.2). Although heterogeneous nucleation sites can bias droplet position (section 4.1.4), true spatial control might require active catalytic sites (section 5.4), which drive localised droplet formation in an undersaturated system. The timing of droplet formation can be controlled by actively driven bulk reactions, which can suppress phase separation (section 5.2.1). This process requires a continuous expenditure of fuel such as ATP, but has the advantage that droplets form immediately when fuel ceases, thus providing a robust stress response [502].
@@
Droplets need to be in the right place to fulfil their function. To control position, cells can use wetting (section 4) to anchor droplets at particular structures, such as the plasma membrane or chromosome loci, so that they remain in place. If droplets are not anchored, cells can reposition them, e.g., using external gradients (section 2.3.2), hydrodynamic fluxes, or active transport [508]. Even more exciting are droplets that swim themselves, either because they self-organise chemical gradients (section 5.5.2) or they exert active stresses [89, 509, 510].
@@ 6.2 Droplet–environment interactions
The interplay of droplets with each other and their environment leads to complex interaction networks, which can process information in cells and enable communication between droplets [516, 517]. Such interaction networks typically involve multiple feedback loops, which give rise to additional dynamic behaviours [518]. At an abstract level, droplets might thus play an integral part in information processing and learning in biological cells [348, 519, 520, 521].
@@ 6.3 Outlook
Beyond exploring how droplets enable cellular function, understanding the physical principles that biology exploits so effectively has application beyond cell biology. For example, it has been hypothesised that phase separation could be implicated in the origin of life by enabling the formation of basic protocells [148]. Indeed, droplets provide well-defined entities that grow, move, and divide by metabolising material from their environment [523, 524]. On the more theoretical side, chemically active droplets comprise an exciting class of non-equilibrium systems [89], with similarities to proliferating active matter [525] and general dissipative structures [526]. Finally, understanding droplet regulation in biological cells will allow us to leverage similar principles in designing novel synthetic systems [65, 70].
We would like to thank Jan Kirschbaum and Estefania Vidal-Henriquez for early input and drafting initial concepts for the review. We thank Kristian Blom, Rumiana Dimova, Marcel Ernst, Stefan Köstler, Guido Kusters, Chengjie Luo, Yicheng Qiang, Riccardo Rossetto, Henri Schmidt, Filipe Thewes, Gerrit Wellecke, Ned Wingreen, Mengmeng Wu, and Noah Ziethen for helpful comments and discussions of the manuscript. We further thank the entire Zwicker group for invaluable scientific discussions. We gratefully acknowledge funding from the Max Planck Society and the European Union (ERC, EmulSim, 101044662).
@@ A Compressible multi-component fluid
The particle fluxes $\boldsymbol{\hat{J}}_{i}$ can be split into a centre-of-mass motion and relative particle fluxes $\boldsymbol{\hat{j}}_{i}=\boldsymbol{\hat{J}}_{i}-c_{i}\boldsymbol{v}$} [101], implying the constraint
$\sum_{i=0}^{{N_{\mathrm{c}}}}m_{i}\boldsymbol{\hat{j}}_{i}=\boldsymbol{0}\;.$ (A.5)
The dynamics of the flow field $\boldsymbol{v}$ follow from momentum conservation, $\partial_{t}(\rho\boldsymbol{v})=\nabla\cdot\boldsymbol{\sigma}^{\mathrm{tot}}$, where the total stress $\boldsymbol{\sigma}^{\mathrm{tot}}=\boldsymbol{\sigma}-\rho\boldsymbol{v}\boldsymbol{v}$ includes the stress $\boldsymbol{\sigma}$ in the co-moving reference frame [121]. The relative fluxes $\boldsymbol{\hat{j}}_{i}$ can be expanded assuming a linear response to thermodynamic forces [101],
$\boldsymbol{\hat{j}}_{i}$ $=-\sum_{j=0}^{{N_{\mathrm{c}}}}L_{ij}\nabla\bigl{(}\hat{\mu}_{j}+\zeta\bigr{)}-\boldsymbol{\hat{\xi}}_{i}$ for $i=0,\ldots,{N_{\mathrm{c}}}\;,$ (A.6)
where $L_{ij}$ denotes the positive definite mobility matrix, $\zeta$ is a Lagrange multiplier ensuring constraint (A.5), and $\boldsymbol{\hat{\xi}}$ is additive noise that we discuss below.
Since only $N_{\mathrm{c}}$} relative fluxes are independent, we can eliminate the flux of one component, where we arbitrarily choose the solvent ($i=0$). In particular, we can then eliminate $\zeta$ by inserting equation (A.6) into equation (A.5) and solving the deterministic part for $\nabla\zeta$. Inserting the solution into equation (A.6) yields
@@
where we introduced the exchange chemical potentials $\hat{\bar{\mu}}_{j}=\hat{\mu}_{j}-m_{j}m_{0}^{-1}$ relative to solvent and the mobility
@@
which obeys Onsager’s symmetry condition ($\hat{\Lambda}_{ij}=\hat{\Lambda}_{ji}$) and additionally $\sum_{i=0}^{N_{\mathrm{c}}}m_{i}\hat{\Lambda}_{ij}=0$. Finally, we use the fluctuation–dissipation relation [101] to determine the variance of the noise $\boldsymbol{\hat{\xi}}_{i}$ for $i=1,\ldots,{N_{\mathrm{c}}}$,
$\langle\boldsymbol{\hat{\xi}}_{i}(\boldsymbol{r},t)\boldsymbol{\hat{\xi}}_{j}(\boldsymbol{r}^{\prime},t^{\prime})\rangle=2k_{\mathrm{B}}T\hat{\Lambda}_{ij}\mathbb{1}\delta(\boldsymbol{r}-\boldsymbol{r}^{\prime})\delta(t-t^{\prime})\;,$ (A.9)
and the solvent flux reads $\boldsymbol{\hat{j}}_{0}=-m_{0}^{-1}\sum_{i=1}^{N_{\mathrm{c}}}m_{i}\boldsymbol{\hat{j}}_{i}$.
To describe incompressible liquids, we assume a large bulk modulus ($K\rightarrow\infty$), which enforces a constant volume density, $\sum_{i}\nu_{i}c_{i}=1$. In the simple case where all components have equal mass density $m_{i}/\nu_{i}$, the total mass density $\rho$ is then also conserved, implying incompressible flows ($\nabla.\boldsymbol{v}=0$) enforced by hydrostatic pressure. The system state is then conveniently expressed in terms of volume fractions $\phi_{i}=\nu_{i}c_{i}$, with the solvent fraction $\phi_{0}=1-\sum_{i=1}^{N_{\mathrm{c}}}\phi_{i}$. The kinetics described by equations (2.8)–(2.11) follow directly by substituting $\Lambda_{ij}=\nu_{i}\hat{\Lambda}_{ij}$ and $\boldsymbol{j}_{i}=\nu_{i}\boldsymbol{\hat{j}_{i}}$. The associated mobility matrix given by equation (2.12) follows directly from equation (A.8). In the simple case of a non-interacting system ($h=0$), diagonal mobilities, $L_{ij}=\lambda_{i}\phi_{i}\delta_{ij}$, dilute concentrations ($\phi_{i}\ll 1$ for $i\geq 1$), and a fast solvent ($\lambda_{0}\gg\lambda_{i}$), the kinetics reduce to ideal diffusion, $\boldsymbol{j}_{i}=-D_{i}\nabla\phi_{i}$, with diffusivity $D_{i}=k_{\mathrm{B}}T\lambda_{i}$.
@@
where $e_{i}$ are internal energies and $e_{ij}$ accounts for pairwise interactions with $e_{ij}=e_{ji}$. This model can be derived from a lattice model or can be viewed as a simple polynomial expansion up to second order, which is sufficient to describe phase separation, although higher-order effects are also interesting [351]. Taken together, this results in the incompressible free energy density given in equation (2.1), where
@@
with $\chi_{ij}=\chi_{ji}$, but $\chi_{ii}\neq 0$. The exchange chemical potentials in equation (2.2) follow from equation (A.3) using $\bar{\mu}_{i}=\hat{\mu}_{i}-\Omega_{i}\hat{\mu}_{0}$ with $\Omega_{i}=\nu_{i}/\nu_{0}$. Similarly, the osmotic pressure $P=\hat{P}-\hat{\mu}_{0}\nu_{0}^{-1}$ in equation (2.3) follows from equation (A.4), illustrating that $P$ involves the chemical potential of the solvent.